OpenStudy (anonymous):

How many positive integers N from 1 to 5000 satisfy the congruence N \equiv 5 \pmod{12}?

OpenStudy (anonymous):

actually, it is N==5(mod 12) the "=="sign means congruent.

OpenStudy (mathmath333):

\(\large \begin{aligned} \color{black}{ N\equiv 5\pmod {12}\\~\\ N=12k+5\\~\\ \normalsize \text{for integers between 1 and 5000}\\ \implies 1\leq 12k+5\leq 5000 }\end{aligned}\)

OpenStudy (anonymous):

I think it will simplify to -4<12k<4995. which also simplifies to 1<k<416. total 417... right! I got it!

OpenStudy (mathmath333):

was it 417

OpenStudy (mathmath333):

i m getting \(416\)

OpenStudy (mathmath333):

or may be i m wrong